OpenStudy (anonymous):

If sin(theta) = -0.89 determine two possible values for theta if 0 degrees < theta < 360 degrees.

OpenStudy (tanya123):

θ, θ, 0, 360 Degree

OpenStudy (campbell_st):

well sin is negative in the 3rd and 4th quadrants so find the value of theta \[\theta = \sin^{-1}(0.89)\] then the 3rd quadrant angle is \[180 + \theta\] and the 4th quadrant angle is \[360 - \theta\]